# Two-Edge Connectivity via Pac-Man Gluing

### v1

@@ 6.3.1 Proof of Lemma 12
[…]
Case 1.2.2: Since we are in Case 1.2 and not in Case 1.2.1, there only exist paths $P_{1},P_{2}$ and $w_{1},w_{2}\in\{v_{1},v_{2},v_{3}\}$ such that $u_{1}\neq u_{2}$ and $w_{1}\neq w_{2}$. Since there is no Hamiltonian path between $u_{1}$ and $u_{2}$ (Case 1.1), it must be that $u_{1}$ and $u_{2}$ are not adjacent in the cycle $C_{A}$. Additionally, $C_{A}$ must be a $\mathcal{C}_{5}$ ($C_{A}=\mathcal{C}_{4}$, $u_{1}\neq u_{2}$, and $u_{1}$ not adjacent to $u_{2}$ implies they are diagonally opposite, and we can set $w_{1}=w_{2}\in\{v_{1},v_{2},v_{3}\}$). Furthermore, we can set $w_{1}$ and $w_{2}$ such that they are adjacent and in $\{v_{1},v_{2},v_{3}\}$ (this can be seen from the case we are in: $u_{1}\neq u_{2}$, $u_{1}$ and $u_{2}$ are not adjacent, $C_{A}=\mathcal{C}_{5}$ and it is not possible to set $w_{1}=w_{2}\in\{v_{1},v_{2},v_{3}\}$). Let $A_{i}\neq A$ be the node in $S$ that is adjacent to $w_{i}$, for $i\in\{1,2\}$. Since $S$ is 2-node connected, there must be a path $P^{\prime}_{i}$ in $S$ from $A_{i}$ to $L$ which does not contain $A$, for $i\in\{1,2\}$.
[…]
Case 2: The vertices $v_{1}$, $v_{3}$, and $v_{4}$ have incident edges to nodes in $S$ other than $A$. As discussed above, this case can only apply if $A$ corresponds to a $\mathcal{C}_{5}$. Further, note that neither $v_{2}$ nor $v_{5}$ can have an incident edge to a node in $S$ other than $A$, as otherwise we are also in Case 1 by simply relabelling the $v_{i}$’s cyclically so that $v_{1},v_{2},v_{3}$ have such incident edges. We have two subcases based on whether or not $v_{2}$ or $v_{5}$ has an incident edge to a component outside of segment $S$.
[…]
Case 2.2: There is an edge between $v_{2}$ (or $v_{5}$ by symmetry) and a node $D$ of ^$\bm{\hat{}}{G}_{H}$ which is a trivial segment $S^{\prime}$ or lies in a non-trivial segment $S^{\prime}$, and $S^{\prime}\neq S$. Note that $A$ must be a cut node in ^$\bm{\hat{}}{G}_{H}$ in this case. Observe that $|V(S)|-|V(C_{A})|\geq 10$, because $S$ contains the node $L$ ($C_{L}$ is huge with 10 vertices) other than $A$. If $|V(S^{\prime})\cup V(C_{A})|\geq 10$, Lemma 3 guarantees the existence of a $4$-matching $M_{4}$ between $V(S)\setminus V(C_{A})$ and $V(C_{A})$ (as $C_{A}$ is a separator, $V(G)$ can be partitioned in 2 parts with $V(C_{A})\cup V(S^{\prime})$ on one side and $V(S)\setminus V(C_{A})$ on the other, so that the only edges crossing are between $V(C_{A})$ and $V(S)\setminus V(C_{A})$). But this means that we are actually in Case 1 as some three consecutive vertices of $C_{A}$ must be incident to the other nodes of $S$.
[…]
Appendix A Reduction to Structured Graphs
This section is dedicated to the proof of Definition 5, which we first restate.
[…]
We assume throughout this section that $\alpha\geq\frac{5}{4}$ and $\varepsilon\in(0,\frac{1}{24}]$. Let $\mathsf{ALG}$ be a polynomial-time algorithm that, given an $(\alpha,\varepsilon)$-structured graph $G^{\prime}=(V^{\prime},E^{\prime})$, returns an $\alpha$-approximate 2EC spanning subgraph $\mathsf{ALG}(G^{\prime})\subseteq E^{\prime}$ of $G^{\prime}$. In the following, we define a reduction algorithm which recursively modifies and partitions the input graph $G$ until we obtain $(\alpha,\varepsilon)$-structured graphs, then solves these with $\mathsf{ALG}$, and finally puts the resulting solutions together to an overall solution for $G$. In particular, our reduction checks for structures that are forbidden in $(\alpha,\varepsilon)$-structured graphs. These are graphs of constant size (w.r.t. $\varepsilon$), parallel edges and self loops, 1-vertex cuts, small $\alpha$-contractible subgraphs (w.r.t. $\varepsilon$), irrelevant edges, non-isolating 2-vertex cuts and large 3-vertex cuts (cf. Definition 5). For all structures except large 3-vertex cuts we can use the reduction presented in [18]. The main contribution of this section is to show how to reduce to graphs without large 3-vertex cuts, given that all other substructures have already been removed.
@@ A.1 Definitions and the Algorithm
Next, we classify an optimal solution $\mathsf{OPT}(G)$ with respect to a large 3-vertex cut.
Definition 13 (Solution types).
@@
$H^{\prime}$ is a $C_{1}-C_{k}$ path of length at least $1$ such that $u,v,w\in C_{1}\cup C_{k}$ and |$\lvert\{u,v,w\}\cap C_{i}\rvert\leq 2$ for $i\in\{1,k\}$. We assume w.l.o.g. that $C_{1}=C(u)=C(v)$.
@@
$H^{\prime}$ is composed of two isolated super-nodes $C_{1}$ and $C_{2}$, where |$\lvert\{u,v,w\}\cap C_{i}\rvert\leq 2$ for $i\in\{1,2\}$. We assume w.l.o.g. that $C_{1}=C(u)=C(v)$.
@@
Let $\mathcal{T}\coloneq\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\}$ be the set of 3-vertex cut optimal solution types.
@@
The following proposition lists compatible solution types of both sides of a 3-vertex cut. The proof is straightforward and, thus, omitted.
@@
Let $\{u,v,w\}$ be a large 3-vertex cut in a graph $G$, and $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v,w\}$ such that $7\leq\lvert V_{1}\rvert\leq\lvert V_{2}\rvert$. Let $G_{1}=G[V_{1}\cup\{u,v,w\}]$, $G_{2}=G[V_{1}\cup\{u,v,w\}]\setminus\{uv,vw,uw\}$ and, for a fixed optimal solution $\mathsf{OPT}(G)$, let $\mathsf{OPT}_{i}=\mathsf{OPT}(G)\cap E(G_{i})$ for $i\in\{1,2\}$. Then, with respect to $\{u,v,w\}$, if
@@
    $\mathsf{OPT}_{1}$ is of type $\mathsf{A}$, $\mathsf{OPT}_{2}$ must be of type $t\in\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\}=\mathcal{T}$.
@@
We are now ready to state our main algorithm $\mathsf{RED}$ (Algorithm 1). The reductions applied in Algorithms 1, 1, 1, 1 and 1 are identical to the reductions given in [18]. When Algorithm 1 reaches Algorithm 1 it calls Algorithm 2 to handle a large 3-vertex cut $\{u,v,w\}$. Whenever the reduction returns edges $K$ of contracted graphs $G|H$, i.e., $K=\mathsf{RED}(G|H)$, we interpret these as the corresponding edges after decontracting $H$, e.g., in Algorithm 1 in Algorithm 1 or in Algorithm 2 in Algorithm 2.
Hence, from now on let us focus on Algorithm 2. In Section A.2 we prove that the claimed sets of edges $F$ in Algorithm 2 exist and are of constant size. The computations done in Algorithm 2 can be done as explained in [18]. That is, we first compute via enumeration minimum-size subgraphs $\overline{\mathsf{OPT}}_{1}^{t}$ of $G_{1}$ of every type $t\in\mathcal{T}$ if such a solution exists, because $G_{1}$ is of constant size in this case. If $\overline{\mathsf{OPT}}_{1}^{t}$ does not exist, $\mathsf{OPT}_{1}^{t}$ also cannot exist. Otherwise, that is, $\overline{\mathsf{OPT}}_{1}^{t}$ exists, we check whether $G_{2}$ is a graph that admits a solution that is compatible with $\overline{\mathsf{OPT}}_{1}^{t}$ according to Proposition 18, and, if that is the case, we set $\mathsf{OPT}_{1}^{t}\coloneq\overline{\mathsf{OPT}}_{1}^{t}$. Finally note that if the algorithm reaches Algorithm 1 in Algorithm 1, then $G$ must be $(\alpha,\varepsilon)$-structured.
@@
1 if $V(G)\leq\frac{4}{\varepsilon}$ then
@@
10if $G$ contains an $\alpha$-contractible subgraph $H$ with $V(H)\leq\frac{4}{\varepsilon}$  then
@@
Input: A 2EC graph $G=(V,E)$ that does not contain cut nodes, non-isolating 2-vertex cuts, or $\alpha$-contractible subgraphs with at most $\frac{4}{\varepsilon}$ vertices. A large 3-vertex cut $\{u,v,w\}$ in $G$.
@@
9       Let $\mathsf{OPT}_{1}^{t}$ be the minimum-size subgraphs of $G_{1}$ for every type $t\in\mathcal{T}$ (w.r.t. $\{u,v,w\}$ and $G_{1}$) that belong to some 2EC spanning subgraph of $G$, if exists. Let $\mathsf{OPT}_{1}^{\min}$ be the existing subgraph of minimum size among all types in $\mathcal{T}$, where ties are broken according to Definition 14 (that is, $t^{\prime}\in\mathcal{T}$ is preferred over $t\in\mathcal{T}$ if $t^{\prime}\succ t$), and $t^{\min}\in\mathcal{T}$ the corresponding type. Let $\mathsf{opt}_{1}^{t}\coloneq|\mathsf{OPT}_{1}^{t}|$ for every $t\in\mathcal{T}$ and $\mathsf{opt}_{1}^{\min}\coloneq|\mathsf{OPT}_{1}^{\min}|$.
@@
22      else if $t^{\min}=\mathsf{C2}$ then
23             if Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains a $C(u)-C(v)$ path. then
@@
28            else if Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains either a $C(u)-C(v)$ or a $C(v)-C(w)$ path. then
@@
A.2 Helper Lemmas

### v2

@@ Appendix A Reduction to Structured Graphs
We assume throughout this section that $\alpha\geq\frac{5}{4}$ and $\varepsilon\in(0,\frac{1}{24}]$. Let $\mathsf{ALG}$ be a polynomial-time algorithm that, given an $(\alpha,\varepsilon)$-structured graph $G^{\prime}=(V^{\prime},E^{\prime})$, returns an $\alpha$-approximate 2EC spanning subgraph $\mathsf{ALG}(G^{\prime})\subseteq E^{\prime}$ of $G^{\prime}$. In the following, we define a reduction algorithm which recursively modifies and partitions the input graph $G$ until we obtain $(\alpha,\varepsilon)$-structured graphs, then solves these with $\mathsf{ALG}$, and finally puts the resulting solutions together to an overall solution for $G$. In particular, our reduction checks for structures that are forbidden in $(\alpha,\varepsilon)$-structured graphs. These are graphs of constant size (w.r.t. $\varepsilon$), parallel edges and self loops, 1-vertex cuts, small $\alpha$-contractible subgraphs (w.r.t. $\varepsilon$), irrelevant edges, non-isolating 2-vertex cuts and large 3-vertex cuts (cf. Definition˜5). For all structures except large 3-vertex cuts, we can use the reduction presented in [GGA23]. The main contribution of this section is to show how to reduce to graphs without large 3-vertex cuts, given that all other substructures have already been removed.
This section is structured as follows. In Section˜A.1 we first introduce definitions to classify solutions regarding 2-vertex cuts and 3-vertex cuts, and then give the formal description of the reduction algorithm, which is split into three algorithm environments, where Algorithm˜1 is the main algorithm, which calls Algorithms˜2 and 3 as subroutines. In Section˜A.2, we give auxiliary results, mainly regarding the removal of 2-vertex cuts and 3-vertex cuts. Finally, in Section˜A.3 we combine these auxiliary results together and give the main proof of Definition˜5.
@@ A.1 Definitions and the Algorithm
A.1.1 Definitions for Non-Isolating 2-Vertex Cuts
Let $\{u,v\}$ be a $2$-vertex cut in a graph $G$. Note that that we can find a partition $(V_{1},V_{2})$ of $V\setminus\{u,v\}$ such that $V_{1}\neq\emptyset\neq V_{2}$ and there are no edges between $V_{1}$ and $V_{2}$. Once we delete $\{u,v\}$ from $G$, the graph breaks into various components such that each is either part of $V_{1}$ or $V_{2}$. Similarly, $\mathsf{OPT}(G)$ breaks into various components once we remove $\{u,v\}$ such that each component is either part of $V_{1}$ or $V_{2}$. In the following definition, we consider different types of graphs that match $H\coloneq\mathsf{OPT}(G)[V_{1}\cup\{u,v\}]$ (or analogously $\mathsf{OPT}(G)[V_{2}\cup\{u,v\}]$).
Definition 13 (2-vertex cut solution types).
Let $H$ be a graph and $u,v\in V(H)$. Let $H^{\prime}$ be obtained by contracting each 2EC component $C$ of $H$ into a single super-node $C$, and let $C(x)$ be the corresponding super-node of the 2EC component of $H$ that contains $x\in V(H)$. We define the following types of $H^{\prime}$ with respect to $\{u,v\}$:
Type $\mathsf{A}$:
$H^{\prime}$ is composed of a single super-node $H^{\prime}=C(u)=C(v)$.
Type $\mathsf{B}$:
$H^{\prime}$ is a C(u)C(v)C(u)-C(v) path of length at least $1$.
Type $\mathsf{C}$:
$H^{\prime}$ consists of two isolated super-nodes $C(u)$ and $C(v)$.
Let $\mathcal{T}_{2}\coloneq\{\mathsf{A},\mathsf{B},\mathsf{C}\}$ be the set of 2-vertex cut optimal solution types.
Definition 14 (Order on types).
$\mathsf{A}\succ\mathsf{B}\succ\mathsf{C}$.
A.1.2 Definitions for Large 3-Vertex Cuts
Similarly to 2-vertex cuts and Definition˜13, we give in the following definition types that match the subgraph of an optimal solution that is induced by the partition corresponding to a large $3$-vertex cut $\{u,v,w\}$ in $G$. Naturally, here are more patterns that can occur compared to 2-vertex cuts.
Definition 15 (3-vertex cut solution types).
@@ A.1.2 Definitions for Large 3-Vertex Cuts
$H^{\prime}$ is a $C_{1}-C_{k}$ path of length at least $1$ such that $u,v,w\in C_{1}\cup C_{k}$ and $\lvert\{u,v,w\}\cap C_{i}\rvert\leq 2$ for $i\in\{1,k\}$. We assume w.l.o.g. that $C_{1}=C(u)=C(v)$.
@@
$H^{\prime}$ is composed of two isolated super-nodes $C_{1}$ and $C_{2}$, where $\lvert\{u,v,w\}\cap C_{i}\rvert\leq 2$ for $i\in\{1,2\}$. We assume w.l.o.g. that $C_{1}=C(u)=C(v)$.
@@
Let $\mathcal{T}_{3}\coloneq\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\}$ be the set of 3-vertex cut optimal solution types.
@@
The following proposition lists compatible solution types of both sides of a 3-vertex cut $\{u,v,w\}$. The proof is straightforward and, thus, omitted. For example, if $\mathsf{OPT}_{1}$ (using the notation introduced below) is of type $\mathsf{C3}$, it is composed of three distinct 2EC components $C(u)$, $C(v)$, and $C(w)$. Therefore, $u$, $v$, and $w$ must be in the same 2EC connected component in $\mathsf{OPT}_{2}$ as otherwise $\mathsf{OPT}$ cannot be a 2-ECSS, and thus, $\mathsf{OPT}_{2}$ must be of type $\mathsf{A}$.
@@
Let $\{u,v,w\}$ be a large 3-vertex cut in a graph $G$, and $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v,w\}$ such that $7\leq\lvert V_{1}\rvert\leq\lvert V_{2}\rvert$ and there are no edges between $V_{1}$ and $V_{2}$ in $G$. Let $G_{1}=G[V_{1}\cup\{u,v,w\}]$, $G_{2}=G[V_{2}\cup\{u,v,w\}]\setminus\{uv,vw,uw\}$ and, for a fixed optimal solution $\mathsf{OPT}(G)$, let $\mathsf{OPT}_{i}=\mathsf{OPT}(G)\cap E(G_{i})$ for $i\in\{1,2\}$. Then, with respect to $\{u,v,w\}$, if
@@
    $\mathsf{OPT}_{1}$ is of type $\mathsf{A}$, $\mathsf{OPT}_{2}$ must be of type $t\in\{\mathsf{A},\mathsf{B1},\mathsf{B2},\mathsf{C1},\mathsf{C2},\mathsf{C3}\}=\mathcal{T}_{3}$.
@@
A.1.3 The Full Algorithm
We are now ready to state our main algorithm $\mathsf{RED}$ (Algorithm˜1). Let $\mathsf{red}(G)\coloneq|\mathsf{RED}(G)|$ denote the size of the solution computed by $\mathsf{RED}$ for a given 2-ECSS instance $G$. The reductions applied in Algorithms˜1, 1, 1, 1 and 1 (here and in the following, we do not distinguish between the line where the actual reduction is computed and the line in which the condition for the reduction is checked) are identical to the reductions given in [GGA23]. Specifically, Algorithm˜2, which is called in Algorithm˜1, reduces non-isolating 2-vertex cuts and was first given in [GGA23]. We emphasize that we present it slightly different, but without changing its actual behavior. When Algorithm˜1 reaches Algorithm˜1 it calls Algorithm˜3 to handle a large 3-vertex cut $\{u,v,w\}$.
Whenever the reduction returns edges $K$ of contracted graphs $G|H$, i.e., $K=\mathsf{RED}(G|H)$, we interpret these as the corresponding edges after decontracting $H$, e.g., in Algorithm˜1 in Algorithm˜1 or in Algorithm˜3 in Algorithm˜3.
Hence, from now on let us focus on Algorithm˜3. In Section˜A.2.2 we prove that the claimed sets of edges $F$ in Algorithm˜3 exist and are of constant size. The computations done in Algorithm˜3 (and in Algorithm˜2 in Algorithm˜2) can be done as explained in [GGA23]. That is, we first compute via enumeration minimum-size subgraphs $\overline{\mathsf{OPT}}_{1}^{t}$ of $G_{1}$ of every type $t\in\mathcal{T}_{3}$ if such a solution exists, because $G_{1}$ is of constant size in this case. If $\overline{\mathsf{OPT}}_{1}^{t}$ does not exist, $\mathsf{OPT}_{1}^{t}$ also cannot exist. Otherwise, that is, $\overline{\mathsf{OPT}}_{1}^{t}$ exists, we check whether $G_{2}$ is a graph that admits a solution that is compatible with $\overline{\mathsf{OPT}}_{1}^{t}$ according to Proposition˜17, and, if that is the case, we set $\mathsf{OPT}_{1}^{t}\coloneq\overline{\mathsf{OPT}}_{1}^{t}$.
Finally, note that if the algorithm reaches Algorithm˜1 in Algorithm˜1, then $G$ must be $(\alpha,\varepsilon)$-structured.
@@ A.1.3 The Full Algorithm
1 if $|V(G)|\leq\frac{4}{\varepsilon}$ then
@@
10if $G$ contains an $\alpha$-contractible subgraph $H$ with $|V(H)|\leq\frac{4}{\varepsilon}$ then
@@
Input: A 2EC graph $G=(V,E)$ without cut vertices and without $\alpha$-contractible subgraphs with at most $\frac{4}{\varepsilon}$ vertices. A non-isolating 2-vertex cut $\{u,v\}$ in $G$.
1 Let $(V_{1},V_{2})$ be a partition of $V\setminus\{u,v\}$ such that $2\leq\lvert V_{1}\rvert\leq\lvert V_{2}\rvert$ and there are no edges between $V_{1}$ and $V_{2}$ in $G$.
2 Let $G_{1}\coloneq G[V_{1}\cup\{u,v\}]$ and $G_{2}\coloneq G[V_{2}\cup\{u,v\}]\setminus\{uv\}$.
3 Let $G^{\prime}_{i}\coloneq G_{i}|\{u,v\}$ for $i\in\{1,2\}$.
4 if $\lvert V_{1}\rvert>\frac{2}{\varepsilon}-4$ then
5  Let $H^{\prime}_{i}\coloneq\mathsf{RED}(G^{\prime}_{i})$ for $i\in\{1,2\}$.
7
8else
9  Let $\mathsf{OPT}_{1}^{t}$ be the minimum-size subgraphs of $G_{1}$ for every type $t\in\mathcal{T}_{2}$ (w.r.t. $\{u,v\}$ and $G_{1}$) that belong to some 2EC spanning subgraph of $G$, if exists. Let $\mathsf{OPT}_{1}^{\min}$ be the existing subgraph of minimum size among all types in $\mathcal{T}_{2}$, where ties are broken according to Definition˜14 (that is, $t^{\prime}\in\mathcal{T}_{2}$ is preferred over $t\in\mathcal{T}_{2}$ if $t^{\prime}\succ t$), and $t^{\min}\in\mathcal{T}_{2}$ the corresponding type. Let $\mathsf{opt}_{1}^{t}\coloneq|\mathsf{OPT}_{1}^{t}|$ for every $t\in\mathcal{T}_{2}$ and $\mathsf{opt}_{1}^{\min}\coloneq|\mathsf{OPT}_{1}^{\min}|$.
10  if $t^{\min}=\mathsf{B}$ then
11     Let $G_{2}^{\mathsf{B}}\coloneq(V(G_{2})\cup\{w\},E(G_{2})\cup\{uw,vw\})$ and $H_{2}^{\mathsf{B}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{B}})\setminus\{uw,vw\}$, where $w$ is a dummy vertex and $uw,vw$ are dummy edges.
12     $H^{\mathsf{B}}\coloneq\mathsf{OPT}_{1}^{\mathsf{B}}\cup H_{2}^{\mathsf{B}}$ return $H^{\mathsf{B}}$.
13
14 else if $t^{\min}=\mathsf{C}$ then
15     Let $G_{2}^{\mathsf{C}}\coloneq(V(G_{2}),E(G_{2})\cup\{uv\})$ and $H_{2}^{\mathsf{C}}\coloneq\mathsf{RED}(G_{2}^{\mathsf{C}})\setminus\{uv\}$.
16     Let $F^{\mathsf{C}}\subseteq E$ be a min-size edge set s.t. $H^{\mathsf{C}}\coloneq\mathsf{OPT}_{1}^{\mathsf{C}}\cup H_{2}^{\mathsf{C}}\cup F^{\mathsf{C}}$ is 2EC. return $H^{\mathsf{C}}$.
17
18
Algorithm 2 Remove a non-isolating 2-vertex cut [GGA23]
Input: A 2EC graph $G=(V,E)$ without cut vertices, non-isolating 2-vertex cuts, or $\alpha$-contractible subgraphs with at most $\frac{4}{\varepsilon}$ vertices. A large 3-vertex cut $\{u,v,w\}$ in $G$.
@@
9  Let $\mathsf{OPT}_{1}^{t}$ be the minimum-size subgraphs of $G_{1}$ for every type $t\in\mathcal{T}_{3}$ (w.r.t. $\{u,v,w\}$ and $G_{1}$) that belong to some 2EC spanning subgraph of $G$, if exists. Let $\mathsf{OPT}_{1}^{\min}$ be the existing subgraph of minimum size among all types in $\mathcal{T}_{3}$, where ties are broken according to Definition˜16 (that is, $t^{\prime}\in\mathcal{T}_{3}$ is preferred over $t\in\mathcal{T}_{3}$ if $t^{\prime}\succ t$), and $t^{\min}\in\mathcal{T}_{3}$ the corresponding type. Let $\mathsf{opt}_{1}^{t}\coloneq|\mathsf{OPT}_{1}^{t}|$ for every $t\in\mathcal{T}_{3}$ and $\mathsf{opt}_{1}^{\min}\coloneq|\mathsf{OPT}_{1}^{\min}|$.
@@
22 else if $t^{\min}=\mathsf{C2}$ then // from now on $y,z$ are dummy vertices with incident dummy edges
23     if Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains a C(u)C(v)C(u)-C(v) path. then
@@
28    else if Every $\mathsf{OPT}_{1}^{\mathsf{C2}}$ solution contains either a C(u)C(v)C(u)-C(v) or a C(v)C(w)C(v)-C(w) path. then
@@
A.2 Auxiliary Lemmas
The following two lemmas are standard results, and can be found in, e.g., [HVV19, GGA23].
Lemma 18.
Let $G$ be a 2EC graph, $v$ be a cut vertex of $G$, and $(V_{1},V_{2}),V_{1}\neq\emptyset\neq V_{2}$, be a partition of $V\setminus\{v\}$ such that there are no edges between $V_{1}$ and $V_{2}$. Then, $\mathsf{opt}(G)=\mathsf{opt}(G[V_{1}\cup\{v\}])+\mathsf{opt}(G[V_{2}\cup\{v\}])$.
Lemma 19.
Let $G$ be a 2VC multigraph with $|V(G)|\geq 3$. Then, there exists a minimum size 2EC spanning subgraph of $G$ that is simple, that is, it contains no self loops or parallel edges.
Next, we replicate a lemma of [GGA23] that shows that irrelevant edges are irrelevant for obtaining an optimal 2-ECSS.
Lemma 20 (Lemma 2.1 in [GGA23]).
Let $e=uv$ be an irrelevant edge of a 2VC simple graph $G=(V,E)$. Then there exists a minimum-size 2EC spanning subgraph of $G$ not containing $e$.
Proof.
Assume by contradiction that all optimal 2EC spanning subgraphs of $G$ contain $e$, and let $\mathsf{OPT}$ be one such solution. Define $\mathsf{OPT}^{\prime}\coloneq\mathsf{OPT}\setminus\{e\}$. Clearly, $\mathsf{OPT}^{\prime}$ cannot be 2EC as it would contradict the optimality of $\mathsf{OPT}$. Let $(V_{1},V_{2})$ be any partition of $V\setminus\{u,v\}$, $V_{1}\neq\emptyset\neq V_{2}$, such that there are no edges between $V_{1}$ and $V_{2}$, which must exist since $\{u,v\}$ is a 2-vertex cut. Notice that at least one of $\mathsf{OPT}^{\prime}_{1}\coloneq\mathsf{OPT}^{\prime}[V_{1}\cup\{u,v\}]$ and $\mathsf{OPT}^{\prime}_{2}\coloneq\mathsf{OPT}^{\prime}[V_{2}\cup\{u,v\}]$, say $\mathsf{OPT}^{\prime}_{2}$, needs to be connected, as otherwise $\mathsf{OPT}$ would not be 2EC. We also have that $\mathsf{OPT}^{\prime}_{1}$ is disconnected, as otherwise $\mathsf{OPT}^{\prime}$ would be 2EC. More precisely, $\mathsf{OPT}^{\prime}_{1}$ consists of exactly two connected components $\mathsf{OPT}^{\prime}_{1}(u)$ and $\mathsf{OPT}^{\prime}_{1}(v)$ containing $u$ and $v$, respectively. Assume w.l.o.g. that $|V(\mathsf{OPT}^{\prime}_{1}(u))|\geq 2$. Observe that there must exist an edge $f$ between $V(\mathsf{OPT}^{\prime}_{1}(u))$ and $V(\mathsf{OPT}^{\prime}_{1}(v))$. Otherwise, $u$ would be a 1-vertex cut separating $V(\mathsf{OPT}^{\prime}_{1}(u))\setminus\{u\}$ from $V\setminus(V(\mathsf{OPT}^{\prime}_{1}(u))\cup\{u\})$. Thus, $\mathsf{OPT}^{\prime\prime}\coloneq\mathsf{OPT}^{\prime}\cup\{f\}$ is an optimal 2EC spanning subgraph of $G$ not containing $e$, a contradiction. ∎
The next lemma shows that for both, 3-vertex cuts and 2-vertex cuts, type $\mathsf{A}$ solutions are more expensive than the cheapest solution.
Lemma 21.
If Algorithm˜2 (Algorithm˜3) reaches Algorithm˜2 (Algorithm˜3) and $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists, then $\mathsf{opt}_{1}^{\mathsf{A}}\geq\alpha\cdot\mathsf{opt}_{1}^{\min}+1$.
Proof.
If $\mathsf{OPT}_{1}^{\mathsf{A}}$ exists and $\mathsf{opt}_{1}^{\mathsf{A}}\leq\alpha\cdot\mathsf{opt}_{1}^{\min}$ then $\mathsf{OPT}_{1}^{\mathsf{A}}$ is 2EC and $\alpha$-contractible. However, since we assume that we reached Algorithm˜2 (Algorithm˜3), we have that $|V_{1}|\leq\frac{2}{\varepsilon}-4\leq\frac{4}{\varepsilon}$. This is a contradiction, because we would have contracted $G_{1}$ in Algorithm˜1 of Algorithm˜1 earlier. Thus, it must be that $\mathsf{opt}_{1}^{\mathsf{A}}\geq\alpha\cdot\mathsf{opt}_{1}^{\min}+1$. ∎
A.2.1 Auxiliary Lemmas for Algorithm˜2